angelcook2007:

1. (08.01, 8.02 MC) An expression is shown below: 6x2y − 3xy − 24xy2 + 12y2 Part A: Rewrite the expression so the GCF is factored completely. (4 points) Part B: Rewrite the expression completely factored. Show the steps of your work.

VIBE:

3y(4x^2 - x-8xy + 4y) Bit by bit clarification: This issue annoys factorization. The GCF, most prominent normal factor is the Term That has the biggest factor that can isolate through every one of the terms in the polynomial In the articulation 6x^2y − 3xy − 24xy^2 + 12y^2 We can see that the terms 3 and y are regular all through the conditions of the polynomial How about we factor 3y out we have 3y(4x^2 - x-8xy + 4y)

angelcook2007:

So, is that the answer or do I have to factor it out more?

VIBE:

That is the answers.

angelcook2007:

oh okay

angelcook2007:

which one is the complete factorization and which one is the gcf factorization?

VIBE:

the first number is the greatest common factor the second one is the complete factorization.

angelcook2007:

Okay, so 6x^2y − 3xy − 24xy^2 + 12y^2 is gcf or complete factorization?

VIBE:

gcf

angelcook2007:

alr, thanks

VIBE:

np

angelcook2007:

can you help me with this one rq if you don't mind?

VIBE:

sure

angelcook2007:

Part A: The area of a square is (9a2 − 24a + 16) square units. Determine the length of each side of the square by factoring the area expression completely. Show your work. (5 points) Part B: The area of a rectangle is (25a2 − 36b2) square units. Determine the dimensions of the rectangle by factoring the area expression completely. Show your work. (5 points)

VIBE:

@VIBE wrote:
Section A) The length of each side of the square is (4a - 3). Part B) The elements of the square shape are (3a + 5b) by (3a - 5b). Bit by bit clarification: Section A) The space of a square is given by: (16a^2-24a+9) Notice that this looks like an ideal square three fold, to be specific: (a-b)^2 =a^2-2ab+b^2 Indeed, in the event that we can see that: 16a^2-24a+9= (4a)^2-2(4a)(3)+(3)^2 At that point we can let a = 4a and b = 3, and factor all things considered: 16a-24+9=(4a-3)^2 The space of a square is given by: A=8^2 So: (4a-3)^2=8^2 Taking the square base of the two sides produces: 8=4a-3 Along these lines, the length of each side of the square is (4a - 3). Recollect that each of the four sides of a square are same. Part B) The space of a square shape is given by: (9a^2-25b^2) Once more, notice that this looks like the type of the distinction of two squares: a^2-b^2=(a+b)(a-b) Along these lines, in the event that we let a = 3a and b = 5b, we can see that: 9a^2-25b^2 =(3a)^2-(5b)^2 Along these lines, we can factor in like manner: 9a^2-25b^2=(3a+5b)(3a-5b) The space of a square shape is given: A=w•l So: A=(3a+5b)(3a-5b) In this way, the components of the square shape are (3a + 5b) by (3a - 5b).

VIBE:

there you go.

angelcook2007:

thanks